OpenStudy (anonymous):

simplify 5 plus i over i

OpenStudy (anonymous):

mulyiply top and bottom by i

OpenStudy (anonymous):

[5+i]/ i (-i/-i) = 5i+1

OpenStudy (anonymous):

1-5i

OpenStudy (anonymous):

\[\frac{5+i}{i}=-5i+1\]The relationship -i+i=1 can be useful here. estudier is correct.